The angle of a triangle is 15 degrees more than the first. The third angle is 23 degrees less than the first. Find the exact measure of all three triangles.
Answer by nerdybill(7384)

The angle of a triangle is 15 degrees more than the first. The third angle is 23 degrees less than the first. Find the exact measure of all three triangles.
.
Let x = first angle
then
x+15 = second angle
x-23 = third angle
.
The sum of the three interior angles of any triangle is 180.
.
x + x+15 + x-23 = 180
3x - 8 = 180
3x = 188
x = 188/3
x = 62 2/3 degrees (first angle)
.
second angle:
x+15 = 77 2/3 degrees (second angle)
.
third angle:
x-23 = 39 2/3 degrees (third angle)